i need help pleas use the quadriatic formula to solve each of the following quadratic equations.
3x^2 - 2x + 1 = 0

Answer by stanbon(75887)

3x^2 - 2x + 1 = 0
Factor as follows:
(3x+1)(x-1)=0
Since the product is zero one of the factors must be zero:
(3x+1)=0 or (x-1)=0
x = -1/3 or x=1
==============
Cheers,
Stan H.